OpenStudy (anonymous):

f(y) = y + y ^2 - 3 Match each value of y with the corresponding value of the function. 1. f(-2) -1 2. f(-4) -3 3. f(0) 17 4. f(2) 9 5. f(4) 3

OpenStudy (anonymous):

justt replace "y" by -2 (for example) ! what will you get ?

OpenStudy (anonymous):

still there @miadeex

OpenStudy (anonymous):

f(-2)=-2+(-2)^2-3 =-2+4-3 =-1